OpenStudy (anonymous):

Jacob invested $300 in a savings account and earned $81 in interest at the end of 9 years. What was the interest rate?

OpenStudy (ranga):

Formula for simple interest is: I = P*r*t where P is the Principal ($300) r is the interest rate in decimal (?) t is the number of years invested (9) and I is the interest earned ($81). Substitute the values into the formula and solve for r.

OpenStudy (ranga):

The "r" you solve will be in decimal. Multiply it by 100 to express it as a percentage.

OpenStudy (anonymous):

The answer is 3%

OpenStudy (ranga):

3% is the correct answer. :)

OpenStudy (wolf1728):

The formula for finding the rate is more complicated than what Ranga posted. The formula is located here: http://1728.org/compint2.htm The correct formula for solving the interest rate is: log(1 + rate) = {log(total) -log(Principal)} / Years log(1 + rate) = {log(381) - log(300)} / Years log(1 + rate) = (2.5809249757 - 2.4771212547) / 9 log(1 + rate) = 0.103803721 / 9 log(1 + rate) = 0.0115337468 We must raise 10 to the power of 0.0115337468 to get the value of 1 + rate 1 + rate = 10^0.0115337468 1 + rate = 1.0269132247 rate = .0269132247 rate = 2.69132247 % And here's a calculator for checking that answer: http://1728.org/compint.htm

OpenStudy (anonymous):

lol ... Ranga's post is correct for the level involved.